1) Consider the following equilibrium: Fe3+(aq) + SCN-(aq) ⇌ FeSCN2+(aq) Initial concentrations: [Fe3+] = 0.590; [SCN-] = 1.239; [FeSCN2+] = 0 The equilibrium concentration of [FeSCN2+]eq = 0.454 M. What is the numerical value of KC for this equilibrium? KC = __________________

2) Consider the following equilibrium: Fe3+(aq) + SCN-(aq) ⇌ FeSCN2+(aq) Initial concentrations: [Fe3+] = 0.370; [SCN-] = 0.777; [FeSCN2+] = 0 The equilibrium concentration of [FeSCN2+]eq = 0.285 M. What is the equilibrium concentration of Fe3+? [Fe3+]eq = _____________M

3) When the kinetic data is correctly plotted to determine the activation energy, Ea, for a chemical reaction, the slope for the straight line graph is found to be -4900. What is the activation energy, Ea, in kJ/mol? Enter your answer as a POSITIVE number. Ea = __________________kJ/mol

4) A standard reference FeSCN2+ solution (0.040 M) has an absorbance = 0.815. What is the concentration of an unkown FeSCN2+ solution that has an absorbance = 0.288. unknown [FeSCN2+] = _________ M

5) 4.00 mL of H2O and 4.00 mL of an aqueous 0.040 M SCN- solution were added to 13.50 mL of an aqueous 0.025 M Fe3+ solution. The total volume of the solution is 21.50 mL. What is the diluted Fe3+ concentration?
Answer #1

1) Consider the following equilibrium: Fe3+(aq) + SCN-(aq) ⇌ FeSCN2+(aq) Initial concentrations: [Fe3+] = 0.590; [SCN-] = 1.239; [FeSCN2+] = 0 The equilibrium concentration of [FeSCN2+]eq = 0.454 M. What is the numerical value of KC for this equilibrium? KC = __________________

Solution:

                              Fe3+(aq) + SCN-(aq) ⇌ FeSCN2+(aq)

Initial                         0.590       1.239           0

Change                      -x              -x             +x

Equilibrium                0.590-x      1.239-x       x

Kc = [FeSCN2+] / [Fe3+] [SCN-] = x / (1.239-x) (0.590-x)

[FeSCN2+] = x = 0.454M , [Fe3+] = 1.239 - 0.454 = 0.785M, [SCN-] = 0.590 - 0.454 = 0.136 M

Kc = 0.454 / 0.785 X 0.136 = 4.25

2) Consider the following equilibrium: Fe3+(aq) + SCN-(aq) ⇌ FeSCN2+(aq) Initial concentrations: [Fe3+] = 0.370; [SCN-] = 0.777; [FeSCN2+] = 0 The equilibrium concentration of [FeSCN2+]eq = 0.285 M. What is the equilibrium concentration of Fe3+? [Fe3+]eq = _____________M

Solution

                    Fe3+(aq) + SCN-(aq) ⇌ FeSCN2+(aq)

Initial             0.370       0.777          0

Change         -x              -x             +x

Equilibrium     0.370-x      0.777-x       x

Kc = [FeSCN2+] / [Fe3+] [SCN-] = x / ( 0.370-x) (0.777-x )

[FeSCN2+] = x = 0.285M , [Fe3+] = 0.370 - 0.285 = 0.085M, [SCN-] = 0.777 - 0.285 = 0.221 M

Kc = 0.285 / 0.085 X 0.221 = 15.17

3) When the kinetic data is correctly plotted to determine the activation energy, Ea, for a chemical reaction, the slope for the straight line graph is found to be -4900. What is the activation energy, Ea, in kJ/mol? Enter your answer as a POSITIVE number. Ea = __________________kJ/mol

solution: In Arrhenius equation , Slope = -Ea/R

Slope = -4900 = -Ea / 8.314 X 10^-3 KJ / mole

Ea = 40.74 KJ / mole

4) A standard reference FeSCN2+ solution (0.040 M) has an absorbance = 0.815. What is the concentration of an unkown FeSCN2+ solution that has an absorbance = 0.288. unknown [FeSCN2+] = _________ M

Solution: The absrobance is directly proportional to concentration

A1 / C1 = A2 / C2

0.815 / 0.04 = 0.288 / [C2]

[C2] = 0.0141 M

5) 4.00 mL of H2O and 4.00 mL of an aqueous 0.040 M SCN- solution were added to 13.50 mL of an aqueous 0.025 M Fe3+ solution. The total volume of the solution is 21.50 mL. What is the diluted Fe3+ concentration?

Solution :

The total volume = 21.50 mL

millimMoles of Fe+3 = Molarity X volume = 0.025 x 13.5

New concentration = 0.025 x 13.5 / 21.50 = 0.0157 M